OpenStudy (anonymous):

24/ ( 2 (3.7 - 1.7))

OpenStudy (anonymous):

24/4=6

OpenStudy (anonymous):

\[24\div(2(3.7-1.7))\] Solve the inner parentheses first;\[3.7-1.7=2\]So,\[24 \div (2(2))\] 2*2=4\[24\div4=6\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You're welcome.